# Performance Trade-off and Joint Waveform Design for MIMO-OFDM DFRC Systems

### v1

@@ III-B Radar-strict problem formulation and the precoding matrix design
[…]
@@
It is assumed that $E\left\{{\boldsymbol{S}(k){\boldsymbol{S}^{H}}(k)}\right\}={E_{s}}{\boldsymbol{I}_{{N_{R}}}},{\rm{}}k=0,1,\ldots,N-1$ Therefore, problem OP2 is equivalent to
[…]
@@
$\left\|{\boldsymbol{H}(k)\boldsymbol{W}(k){\boldsymbol{W}^{H}}(k){\boldsymbol{H}^{H}}(k)\!-\!{\boldsymbol{I}_{{N_{R}}}}}\right\|\!=\!\left\|{\boldsymbol{H}(k)\boldsymbol{W}(k)\!-\!{\boldsymbol{I}_{{N_{R}}}}}\right\|,$ (23)
and the optimization problem OP3 can be relaxed into
[…]
$\quad\quad{\rm{s.t.}}\sum\limits_{k=0}^{N-1}{{\alpha_{k}}=1},$ (24b)
$\quad\quad\quad{\alpha_{k}}\geq 0,k=0,1,\ldots,N-1.$ (24c)
@@
[…]
@@
The optimization problem OP5 is an Orthogonal Procrustes problem (OPP) problem, and it has a globally optimal solution, which is given by [44]
@@
where ${\boldsymbol{I}_{{N_{T}}\times{N_{R}}}}$ is a diagonal matrix and each diagonal element is equal to one, $\widetilde{\boldsymbol{U}}(k)\boldsymbol{\Sigma}(k){\widetilde{\boldsymbol{V}}^{H}}(k)$ is the Singular Value Decomposition (SVD) of ${\boldsymbol{T}^{H}}(k){\boldsymbol{H}^{H}}(k)$ Finally, the suboptimal precoding matrix can be derived as
@@ III-C DFRC performance trade-off problem formulation and precoding matrix design
Note that in the optimization problem OP2, the transmit covariance matrix strictly conforms to the optimal result of the Radar-only problem in Section 3.1. Because of this, the communication performance is greatly degraded, and the interference among different information streams at the receiver is serious. Therefore, we consider a trade-off problem, and extend the traditional work in [23] to the frequency-selective multipath channel, where communication performance (ISI among different data streams) and the radar performance (CRB of radar detection) are balanced.
The suboptimal precoding matrix derived in Section 3.2 is denoted by ${\boldsymbol{W}_{d}}(k)$, and the trade-off problem is formulated as [23]
[…]
@@
[…]
@@ IV Joint waveform design based on CRB minimization and achievable rate maximization
In [23], [30, 31, 32, 33] and [40], different receive antennas belong to different communication users, and each user is equipped with one antenna. In this paper, a single communication user MIMO scenario is considered, and all the receive antennas belong to the same user. However, there is no related integrated signal design method in the considered scenario. In this section, we propose a joint waveform design strategy based on the CRB minimization and achievable rate maximization.
@@ IV-A Communication-only problem formulation
[…]
@@ IV-B Covariance matrix design for DFRC performance trade-off
[…]
@@
[…]
@@
[…]
@@ IV-C Precoding matrix design for the DFRC performance trade-off
Similarly, ${\boldsymbol{R}_{f,2}}(k)$ should satisfy
$\boldsymbol{R}_{f,2}(k)=\boldsymbol{W}_{2}(k){\boldsymbol{P}_{S,2}}(k){\boldsymbol{W}^{H}_{2}}(k).$ (44)
Since ${{{P_{\rm{1}}}}\over{{P_{\rm{2}}}}}={{{\rho_{2}}}\over{1-{\rho_{2}}}}$, we can set ${\boldsymbol{P}_{S,2}}(k)$ as
${\boldsymbol{P}_{S,2}}(k)={{1-\rho_{2}}\over\rho_{2}}{\boldsymbol{P}_{S,1}}(k).$ (45)
Define ${{\boldsymbol{P}}^{\circ}}(k)={\left({{\boldsymbol{P}_{S,{\rm{1}}}}(k)}\right)^{1/2}}$, substitute Eq. (45) into Eq. (44), and ${\boldsymbol{R}_{f,{\rm{2}}}}(k)$ can be derived as
${\boldsymbol{R}_{f,{\rm{2}}}}(k)={{1-{\rho_{2}}}\over{{\rho_{2}}}}{\boldsymbol{W}_{\rm{2}}}(k){{\boldsymbol{P}}^{\circ}}(k){\left({{\boldsymbol{W}_{\rm{2}}}(k){{\boldsymbol{P}}^{\circ}}(k)}\right)^{H}}.$ (46)
The sizes of ${\boldsymbol{R}_{f,{\rm{2}}}}(k)$, ${\boldsymbol{W}_{\rm{2}}}(k)$, ${{\boldsymbol{P}}^{\circ}}(k)$ are $N_{T}\times N_{T}$, $N_{T}\times N_{R}$, and $N_{R}\times N_{R}$, respectively. Note that ${{N}_{R}<{N}_{T}}$ is assumed in this paper. Therefore, according to Eq. (46), we can derive the approximate solution of the precoding matrix ${\boldsymbol{W}_{\rm{2}}}(k)$ as [47]
${\boldsymbol{W}_{\rm{2}}}(k)={\boldsymbol{Q}}(k){\left({{\boldsymbol{D}}(k)}\right)^{1/2}}{\boldsymbol{I}_{{N_{T}}\times{N_{R}}}}{\left({{(\boldsymbol{P})^{\circ}}(k)}\right)^{-1}},$ (47)
where ${\boldsymbol{Q}}(k){\boldsymbol{D}}(k)\boldsymbol{Q}^{-1}(k)={{{\rho_{2}}}\over{1-{\rho_{2}}}}{\boldsymbol{R}_{f,{\rm{2}}}}(k)$ is the eigenvalue decomposition of ${\boldsymbol{R}_{f,{\rm{2}}}}(k)$.
Finally, the frequency-domain precoding matrix is designed as
$\boldsymbol{W}(k)=\sqrt{{{1-{\rho_{2}}}\over{{\rho_{2}}}}}{\boldsymbol{W}_{\rm{2}}}(k)+{\boldsymbol{W}_{\rm{1}}}(k),k=0,1,\ldots,N-1.$ (48)
In summary, the waveform design strategy based on the WF-SQP algorithm is provided in Algorithm II.
Algorithm II: The waveform design strategy based on the
@@ 2. For a given $\rho_{2}$, calculate $P_{1}$ and solve the Communication
 problem OP9 through SQP method and obtain $\boldsymbol{\omega}$.
4. Calculate ${\boldsymbol{R}_{f,{\rm{2}}}}(k)$, according to ${\boldsymbol{R}_{f,{2}}}(k)={\omega_{k}}{N^{2}}{\boldsymbol{R}_{2}}$.
5. Calculate the precoding matrix $\boldsymbol{W}_{2}(k)$ based on Eq. (47).
@@
It is noteworthy that PAPR is one of the most important factors in both communication and radar systems, and traditional PAPR reduction strategies [48][49] can be employed in the proposed DFRC system to ensure the effectiveness of radar detection and communication service.
@@ V Simulation results
In this section, we verify the proposed waveform design strategies through simulations. The total transmit power $P_{T}$ is normalized to be 1, and each entry of the frequency-domain channel matrix ${\boldsymbol{H}(k)}$ is subject to standard complex Gaussian distribution. The DFRC BS is equipped with $N_{T}=20$ antennas. The number of receive antennas at the communication user is $N_{R}=10$. The data streams are modulated by quadrature phase shift keying (QPSK). The number of the OFDM subcarriers is $N=16$. The interested detection target is located at the angle of ${0^{\circ}}$. In the figures, the traditional scheme in [23] labelled with ‘Traditional scheme’ is provided as the benchmark scheme, the designs with the strict CRB minimization constraint (mentioned in Section III B) and the performance trade-off are labelled with ‘Strict’ and ‘Trade-off’, respectively, and we use ‘ISI-min’ (inter-stream interference minimization), and ‘AR-max’ (achievable rate maximization) to distinguish the proposed two strategies. Note that the precoding matrix design corresponding to ‘AR-max’ strategy in Eq. (47) is an approximate solution, thus the derived achievable rate using the proposed precoding design will be lower than the ideal optimal result of OP9, and these two cases are labelled as ‘Actual’ and ‘Ideal’, respectively.
The symbol error rates (SERs) of the proposed waveform design strategies are plotted in Fig. 2. Note that the frequency-selective multipath fading channel is considered in this paper, the scheme in [23] labelled by ‘Traditional scheme’ cannot be directly applied to achieve a satisfied result. It can be seen from Fig. 2 that the ‘Strict’ design suffers great communication performance loss under the strict equality constraint (20a)(20b), and the communication performance can be improved by introducing the trade-off factor. Fig. 2 also shows the performance of the proposed ‘Trade-off’ design strategy based on the CRB and ISI minimization, which is an extension of the work in [23] to the frequency-selective multipath fading environment. It can be also seen from Fig. 2 that the second proposed waveform design strategy based on the CRB minimization and the achievable rate maximization achieves the best SER performance.
Figure 2: Symbol error rate comparison of different strategies with different SNR.
Figure 3: Achievable rate comparison of different strategies with different SNR.
Figure 4: Symbol error rate comparison of different strategies with different trade-off factors, SNR=20 dB.
Figure 5: Achievable rate comparison of different strategies with different trade-off factors, SNR=20 dB.
@@
Figure 7: Radar beam pattern obtained by the ‘ISI-min’ strategy with different trade-off factors.
Figure 8: Radar beam pattern obtained by the ‘AR-max’ strategy with different trade-off factors.
Fig. 3 shows the achievable rate of the proposed waveform design strategies. It can be seen from Fig. 3 that the proposed waveform design strategy based on the achievable rate maximization outperforms the strategy based on ISI minimization. Furthermore, it is noteworthy that by employing the proposed precoding matrix design in Eq. (48), the achievable rate only experiences a slight performance loss compared to the ideal optimal result of OP9.
The SER versus the trade-off factor is shown in Fig. 4, where SNR is $20\,\rm{dB}$. It can be seen from Fig. 4 that the SERs of the proposed two waveform design strategies decrease with the increase of the trade-off factor. In the first waveform design strategy (ISI-min), when the trade-off factor is greater than 0.6, there is an error floor. That is because when the first term in Eq. (31a) is close to 0, there is a communication performance bound of this strategy. For the second waveform design strategy (AR-max), there is no error floor, which proves the superiority of the proposed ‘AR-max’ strategy.
Fig. 5 shows achievable rate versus the trade-off factor, and SNR is also equal to 20 dB. The achievable rates of the two waveform design strategies get improved with the increase of trade-off factor. Moreover, the communication performance of the ‘AR-max’ strategy continuously improves with the increase of trade-off factor, while the ‘ISI-min’ based strategy reaches its performance bound when the trade-off factor is greater than 0.6, which is similar to the results in Fig. 4. Meanwhile, the stability of the second proposed waveform design strategy (AR-max) is also verified in Fig. 5, and the gap between the actual result and the ideal optimal result in OP9 is very small.
It is noteworthy that the ‘AR-max’ strategy and ‘ISI-min’ strategy are fundamentally different. The aim of ‘ISI-min’ strategy is to minimize the interference among the signals of different receive antennas, and the signals of different receive antennas do not need to be jointly processed. The aim of the ‘AR-max’ strategy is to maximize the achievable sum rate, and the receiver jointly processes the signals of different receive antennas, which achieves better performance than the ‘ISI-min’ strategy. Therefore, the SER and achievable rate performances of the ‘AR-max’ strategy are better than those of the ‘ISI-min’ strategy.
@@
The resultant radar beampatterns obtained by the two proposed strategies with different trade-off factors are provided in Fig. 7 and Fig. 8, respectively. The ‘Strict’ beam pattern corresponds to the optimal result of OP1. It can be seen from Fig. 7 and Fig. 8 that when the trade-off factor equals 0, the proposed two performance trade-off waveform design strategies obtain the same beam pattern as the ‘Strict’ beam pattern. Meanwhile, with the increase of trade-off factor, the mismatch between the obtained beam pattern and the ‘Strict’ beam pattern begins to increase, because the radar performance is sacrificed to enhance the communication performance.
Figure 9: NMSEs of radar beam pattern obtained by different strategies.
Figure 9: NMSEs of radar beam pattern obtained by different strategies.
Figure 10: Normalized CRBs obtained by different strategies.
@@
where ${B_{S}}(\theta)$ is the ‘Strict’ beam pattern, and ${B_{T}}(\theta)$ is the beam pattern with ‘Trade-off’ strategy.

### v2

@@ III-B Radar-strict problem formulation and the precoding matrix design
$\rm{OP2:}\quad\mathop{\arg\min}\limits_{\boldsymbol{W}(k)}$ $\left({\sum\limits_{k=0}^{N-1}{E\{{{{\left\|{\boldsymbol{H}(k)\boldsymbol{W}(k)\boldsymbol{s}(k)-\boldsymbol{s}(k)}\right\|}^{2}}}\}}}\right)$ (20a)
@@
It is assumed that $E\left\{{\boldsymbol{s}(k){\boldsymbol{s}^{H}}(k)}\right\}={E_{s}}{\boldsymbol{I}_{{N_{R}}}},{\rm{}}k=0,1,\ldots,N-1$ . Therefore, problem OP2 is equivalent to
$\rm{OP3:}\quad$ $\mathop{\arg\min}\limits_{\boldsymbol{W}(k)}\left({\sum\limits_{k=0}^{N-1}{{{\left\|{\boldsymbol{H}(k)\boldsymbol{W}(k)-{\boldsymbol{I}_{{N_{R}}}}}\right\|}^{2}}}}\right)$ (21a)
@@
$\left\|{{\boldsymbol{W}}(k){{\boldsymbol{W}}^{H}}(k)-{{\boldsymbol{H}}^{+}}(k){{\left({{{\boldsymbol{H}}^{+}}(k)}\right)}^{H}}}\right\|=0,$ (23)
where ${{\boldsymbol{H}}^{+}}(k)$ is the Pseudo-inverse of ${\boldsymbol{H}}(k)$. Thus, the optimization problem OP3 can be relaxed into
${{\rm{OP4:}}}$ ${\mathop{{\rm{argmin}}}\limits_{{\boldsymbol{\alpha}}}\sum\limits_{k=1}^{N}{{{\left\|{\frac{{{\alpha_{k}}{N^{2}}}}{{{E_{s}}}}{{\boldsymbol{R}}_{d}}-{{\boldsymbol{H}}^{+}}(k){{\left({{{\boldsymbol{H}}^{+}}(k)}\right)}^{H}}}\right\|}^{2}}}}$ (24a)
${{\rm{s}}{\rm{.t}}{\rm{.}}\sum\limits_{k=1}^{N}{{\alpha_{k}}}=1{}}$ (24b)
${{\rm{}}\quad\quad{\alpha_{k}}\geq 0}$ (24c)
@@
$\rm{OP5:}\quad\mathop{\arg\min}\limits_{\widetilde{\boldsymbol{W}}}$ $\left({{{\left\|{\boldsymbol{H}(k)\boldsymbol{T}(k)\widetilde{\boldsymbol{W}}(k)-{\boldsymbol{I}_{{N_{R}}}}}\right\|}^{2}}}\right)$ (27a)
@@
The optimization problem OP5 is an Orthogonal Procrustes problem (OPP), and it has a globally optimal solution, which is given by [46]
@@
where $\boldsymbol{I}_{{N_{T}}\times{N_{R}}}$} is a diagonal matrix and each diagonal element is equal to one, $\widetilde{\boldsymbol{U}}(k)\boldsymbol{\Sigma}(k){\widetilde{\boldsymbol{V}}^{H}}(k)$ is the Singular Value Decomposition (SVD) of ${\boldsymbol{T}^{H}}(k){\boldsymbol{H}^{H}}(k)$ . Finally, the suboptimal precoding matrix can be derived as
@@ III-C DFRC performance trade-off problem formulation and precoding matrix design
Note that in the optimization problem OP2, the transmit covariance matrix strictly conforms to the optimal result of the Radar-only problem in Section III.A. Because of this, the communication performance is greatly degraded, and the interference among different information streams at the receiver is serious. Therefore, we consider a trade-off problem, and extend the traditional work in [23] to the frequency-selective multipath channel, where communication performance (ISI among different data streams) and the radar performance (CRB of radar detection) are balanced.
The suboptimal precoding matrix derived in Section III.B is denoted by ${\boldsymbol{W}_{d}}(k)$, and the trade-off problem is formulated as [23]
$\rm{OP6:}\quad\mathop{\arg\min}\limits_{\boldsymbol{W}}\sum\limits_{k=0}^{N-1}$ ${{\rho_{1}}{{\left\|{\boldsymbol{H}(k)\boldsymbol{W}(k)\boldsymbol{s}(k)-\boldsymbol{s}(k)}\right\|}^{2}}}$
@@
$\rm{OP7:}\quad$ $\quad\mathop{\arg\min}\limits_{\boldsymbol{W}(k)}{\rho_{1}}{\left\|{\boldsymbol{H}(k)\boldsymbol{W}(k)-{\boldsymbol{I}_{{N_{R}}}}}\right\|^{2}}$
@@ IV Joint waveform design based on CRB minimization and achievable rate maximization
In [23], [32, 33, 34, 35] and [42], different receive antennas belong to different communication users, and each user is equipped with one antenna. In this paper, a single communication user MIMO scenario is considered, and all the receive antennas belong to the same user. However, there is no related integrated signal design method in the considered scenario. In this section, we propose a joint waveform design scheme based on the CRB minimization and achievable rate maximization.
@@ IV-A Communication-only problem formulation
${\rm{OP8:}}\mathop{\arg\max}\limits_{{\boldsymbol{W}}(k),{\boldsymbol{P}_{S}}(k)}\bigg{\{}\sum\limits_{k=0}^{N-1}{{\Delta f}\over B}$
@@ IV-B Covariance matrix design for DFRC performance trade-off
${\rm{OP9:}}\quad\mathop{\arg\max}\limits_{\boldsymbol{\omega}}\sum\limits_{k=0}^{N-1}{{{\Delta f}\over B}}$
@@
where $\boldsymbol{1}=[1,1,\cdots,1]$, and ${\boldsymbol{e}_{k}}=[0,0,\cdots,\underbrace{1}_{k{\rm{th}}},\cdots,0]$ is the unit vector. The optimization problem OP9 can be rewritten by
${\rm{OP10:}}\quad\mathop{\arg\min}\limits_{\omega}f(\boldsymbol{\omega})$ (40a)
@@
${\rm{OP11:}}$ $\quad\mathop{\arg\min}\limits_{\boldsymbol{\eta}}{1\over 2}{\boldsymbol{\eta}^{T}}{\nabla^{2}}f({\boldsymbol{\omega}_{(m)}})\boldsymbol{\eta}+\nabla f{({\boldsymbol{\omega}_{(m)}})^{T}}\boldsymbol{\eta}$ (42a)
@@ IV-C Precoding matrix design for the DFRC performance trade-off
Similarly, ${\boldsymbol{R}_{f}}(k)$ should satisfy
$\boldsymbol{R}_{f}(k)=\boldsymbol{W}(k){\boldsymbol{P}_{S}}(k){\boldsymbol{W}^{H}}(k).$ (44)
For the total transmit power $P_{T}$}, since ${\rho_{2}}={P_{1}}/{P_{T}}$, we can set the optimal power allocation matrix ${{\boldsymbol{P}}_{S}}(k)$ as
${{\boldsymbol{P}}_{S}}(k)=\frac{1}{{{\rho_{2}}}}\cdot{{\boldsymbol{P}}_{S,1}}(k).$ (45)
Define ${\boldsymbol{P}}^{\circ}(k)={({{\boldsymbol{P}}_{S}}(k))^{1/2}}$, and ${{\boldsymbol{R}}_{f}}(k)$ can be derived as
${{\boldsymbol{R}}_{f}}(k)={\boldsymbol{W}}(k){\boldsymbol{P}}^{\circ}(k){\left({{\boldsymbol{W}}(k){\boldsymbol{P}}^{\circ}(k)}\right)^{H}}.$ (46)
The sizes of ${{\boldsymbol{R}}_{f}}(k)$, ${\boldsymbol{W}}(k)$, ${{\boldsymbol{P}}^{\circ}}(k)$ are $N_{T}\times N_{T}$, $N_{T}\times N_{R}$, and $N_{R}\times N_{R}$, respectively. Note that ${N}_{R}<{N}_{T}$} is assumed in this paper. Therefore, through eigenvalue decomposition, we can derive the solution of the precoding matrix ${\boldsymbol{W}}(k)$ as [49]
${\boldsymbol{W}}(k)={\boldsymbol{Q}}(k){\left({{\boldsymbol{D}}(k)}\right)^{1/2}}{\boldsymbol{I}_{{N_{T}}\times{N_{R}}}}{\left({{\boldsymbol{P}^{\circ}}(k)}\right)^{-1}},$ (47)
where ${\boldsymbol{Q}}(k){\boldsymbol{D}}(k)\boldsymbol{Q}^{-1}(k)={\boldsymbol{R}_{f}}(k)$ is the eigenvalue decomposition of ${\boldsymbol{R}_{f}}(k)$.
In summary, the waveform design scheme based on the WF-SQP algorithm is provided in Algorithm II.
Algorithm II: The waveform design scheme based on the
@@ 2. For a given $\rho_{2}$, calculate $P_{1}$ and solve the Communication
   problem OP9 through SQP method and obtain $\boldsymbol{\omega}_{k}$.
4. According to ${\boldsymbol{R}_{f,{2}}}(k)={\omega_{k}}{N^{2}}{\boldsymbol{R}_{2}}$, calculate
${{\boldsymbol{R}}_{f}}(k)={{\boldsymbol{R}}_{f,1}}(k)+{{\boldsymbol{R}}_{f,2}}(k)$.
@@ 4. According to ${\boldsymbol{R}_{f,{2}}}(k)={\omega_{k}}{N^{2}}{\boldsymbol{R}_{2}}$, calculate
It is noteworthy that PAPR is one of the most important factors in both communication and radar systems, and traditional PAPR reduction schemes [50][51] can be employed in the proposed DFRC system to ensure the effectiveness of radar detection and communication service.
@@ V Simulation results
In this section, we verify the proposed waveform design schemes through simulations. The total transmit power $P_{T}$ is normalized to be 1, and each entry of the frequency-domain channel matrix $\boldsymbol{H}(k)$} is subject to standard complex Gaussian distribution. It is noteworthy that more standardized channel model will be adopted in our future work to verify the performance of the proposed scheme. The DFRC BS is equipped with $N_{T}=20$ antennas. The number of receive antennas at the communication user is $N_{R}=10$. The data streams are modulated by quadrature phase shift keying (QPSK). The number of the OFDM subcarriers is $N=16$. The interested detection target is located at the angle of $0^{\circ}$}. In the figures, the traditional scheme in [23] labelled with ‘Traditional scheme’ and the JCAS scheme in [53] labelled with ‘JCAS scheme’ are provided as the benchmark schemes, the designs with the strict CRB minimization constraint (mentioned in Section III B) and the performance trade-off are labelled with ‘Strict’ and ‘Trade-off’, respectively, and we use ‘ISI-min’ (inter-stream interference minimization), and ‘AR-max’ (achievable rate maximization) to distinguish the proposed two schemes. In the simulation, the trade-off factor $\rho$ corresponds to $\rho_{1}$ in the ’ISI-min’ scheme and $\rho_{2}$ in the ’AR-max’ scheme, respectively. Note that the precoding matrix design corresponding to ‘AR-max’ scheme in Eq. (47) is an approximate solution, thus the derived achievable rate using the proposed precoding design will be lower than the ideal optimal result of OP9, and these two cases are labelled as ‘Actual’ and ‘Ideal’, respectively.
The SER of the proposed waveform design schemes are plotted in Fig. 2. Note that the frequency-selective multipath fading channel is considered in this paper, the scheme in [23] labelled by ‘Traditional scheme’ cannot be directly applied to achieve a satisfied result. It can be seen from Fig. 2 that the ‘Strict’ design suffers great communication performance loss under the strict equality constraint (20a)(20b), and the communication performance can be improved by introducing the trade-off factor. Fig. 2 also shows the performance of the proposed ‘Trade-off’ design scheme based on the CRB and ISI minimization, which is an extension of the work in [23] to the frequency-selective multipath fading environment. It can be also seen from Fig. 2 that the second proposed waveform design scheme based on the CRB minimization and the achievable rate maximization achieves the best SER performance.
Figure 2: SER comparison of different schemes with different SNR.
Figure 3: Achievable rate comparison of different schemes with different SNR.
Figure 4: SER comparison of different schemes with different trade-off factors and SNR=20 dB.
Figure 5: Achievable rate comparison of different schemes with different trade-off factors and SNR=20 dB.
@@
Figure 7: Radar beam pattern obtained by the ‘ISI-min’ scheme with different trade-off factors.
Figure 8: Radar beam pattern obtained by the ‘AR-max’ scheme with different trade-off factors.
Fig. 3 shows the achievable rate of the proposed waveform design schemes. It can be seen from Fig. 3 that the proposed waveform design scheme based on the achievable rate maximization outperforms the scheme based on ISI minimization. Furthermore, it is noteworthy that by employing the proposed precoding matrix design in Eq. (47), the achievable rate only experiences a slight performance loss compared to the ideal optimal result of OP9.
The SER versus the trade-off factor is shown in Fig. 4, where SNR is $20\,\rm{dB}$. It can be seen from Fig. 4 that the SERs of the proposed two waveform design schemes decrease with the increase of the trade-off factor. Since the JCAS scheme in [53] is based on the multi-carrier modulation and the scheme in [23] is only suitable for the flat fading channel, the SER of the JCAS scheme is lower than that of the scheme in [23], but higher than that of the proposed two schemes. In the first waveform design scheme (ISI-min), when the trade-off factor is greater than 0.6, there is an error floor. That is because when the first term in Eq. (31a) is close to 0, there is a communication performance bound of this scheme. For the second waveform design scheme (AR-max), there is no error floor, which proves the superiority of the proposed ‘AR-max’ scheme.
Fig. 5 shows achievable rate versus the trade-off factor, and SNR is also equal to 20 dB. The achievable rates of the two waveform design schemes get improved with the increase of trade-off factor. The achievable rate of the JCAS scheme is higher than that of the scheme in [23], but lower than the proposed two schemes. Moreover, the communication performance of the ‘AR-max’ scheme continuously improves with the increase of trade-off factor, while the ‘ISI-min’ based scheme reaches its performance bound when the trade-off factor is greater than 0.6, which is similar to the results in Fig. 4. Meanwhile, the stability of the second proposed waveform design scheme (AR-max) is also verified in Fig. 5, and the gap between the actual result and the ideal optimal result in OP9 is very small.
It is noteworthy that the ‘AR-max’ scheme and ‘ISI-min’ scheme are fundamentally different. The aim of ‘ISI-min’ scheme is to minimize the interference among the signals of different receive antennas, and the signals of different receive antennas do not need to be jointly processed. The aim of the ‘AR-max’ scheme is to maximize the achievable sum rate, and the receiver jointly processes the signals of different receive antennas, which achieves better performance than the ‘ISI-min’ scheme. Therefore, the SER and achievable rate performances of the ‘AR-max’ scheme are better than those of the ‘ISI-min’ scheme.
@@
The resultant radar beampatterns obtained by the two proposed schemes with different trade-off factors are provided in Fig. 7 and Fig. 8, respectively. The ‘Strict’ beam pattern corresponds to the optimal result of OP1. It can be seen from Fig. 7 and Fig. 8 that when the trade-off factor equals 0, the proposed two performance trade-off waveform design schemes obtain the same beam pattern as the ‘Strict’ beam pattern. Meanwhile, with the increase of trade-off factor, the mismatch between the obtained beam pattern and the ‘Strict’ beam pattern begins to increase, because the radar performance is sacrificed to enhance the communication performance.
Figure 9: NMSEs of radar beam pattern obtained by different schemes.
Figure 10: Normalized CRBs obtained by different schemes.
@@
where ${B_{S}}(\theta)$ is the ‘Strict’ beam pattern, and ${B_{T}}(\theta)$ is the beam pattern with ‘Trade-off’ scheme.
@@
Figure 11: Radar beam pattern with different $\theta_{c}$.
Figure 12: SER with different $\theta_{c}$.
Fig. 11 and Fig. 12 show the impact of the angular spacing between the detection target and the communication user on the radar and communication performance. It is assumed that the communication user is located at the angle of $\theta_{c}$ , and the detection target is located at $0^{\circ}$}. By employing the ‘AR-max’ scheme, the transmit beam pattern with different $\theta_{c}$ is shown in Fig. 11. It can be seen from Fig. 11 that when $\rho=0$, that is, the communication performance is not considered, the tramsmit beam radiates highest power to the detection target. When $\rho=0.5$, that is, the radar detection performance and the communication performance are jointly considered, the transmit beam pattern radiates lower power towards the detection target, which will decrease the radar detection performance. Besides, it can be seen from Fig. 11 that the proposed scheme is robust to the angular spacing between the detection target and the communication user. Fig. 12 shows the SER with different $\theta_{c}$. The SER performance of the traditional scheme using single carrier modulation [23] is also provided in Fig. 12 for comparison. It can be seen from Fig. 12 that the proposed waveform design scheme has a better SER performance than the traditional scheme. Fig. 12 also shows that the proposed waveform design scheme is robust to the angular spacing between the detection target and the communication user.